1. What centripetal force is required to cause a 12.0 kg object to complete seven revolutions around a horizontal circle of radius 7.50 m in 24.0 s?


1
Expert's answer
2021-07-22T13:06:09-0400
"F_c=ma_c=\dfrac{mv^2}{r}=\dfrac{m(\omega r)^2}{r}=m\omega^2r,""F_c=12\ kg\cdot(\dfrac{7\ rev}{24\ s}\cdot\dfrac{2\pi\ rad}{1\ rev})^2\cdot7.5\ m=302.3\ N."